OpenStudy (anonymous):

y= - 2/3 X -2 identify the slope and the y-intercept of the equation (I just for got how to do this a bit)

OpenStudy (anonymous):

@bahrom7893

OpenStudy (anonymous):

@dan815

OpenStudy (texaschic101):

y = mx + b.....m is the slope and b is the y-intercept. y = -2/3x - 2.....slope(m) = -2/3 and y-intercept (b) is -2

thomaster (thomaster):

Next time, post questions like these in the \(\sf\color{blue}{\href{ /study#/groups/mathematics}{\underline{mathematics}}}\) section.